OpenStudy (anonymous):

how can i solve the inequality |3x-2| greater than or equal to 5

OpenStudy (tkhunny):

Think about \(|3x-2|\). Where is that positive (or zero) and where is ti negative? This will lead you to a solution.

OpenStudy (anonymous):

you should get 2 answers as well

OpenStudy (tkhunny):

Think about (3x−2). Where is that positive (or zero) and where is ti negative? This will lead you to a solution.

OpenStudy (anonymous):

\[|3x-2|\ge 5\]

OpenStudy (anonymous):

1st thing you need to do is remove the absolute value and set up the 2 equations

OpenStudy (anonymous):

what is the absolute value?

OpenStudy (anonymous):

the lines that the equation is inside of

OpenStudy (anonymous):

oh ok how do i remove that?

OpenStudy (calculusxy):

\[3x-2=5\]

OpenStudy (calculusxy):

Now solve for x. This is the same thing because you may be finding what could be EQUAL to 5.

OpenStudy (anonymous):

\[3x-2 \ge 5 \] and \[ 3x -2 \le -5 \] is what you should get when removing absolute value notice how the 1st equation is exactly the same as the one you wrote down without the absolute value and in the second one you switch the sign and change the number into negative (in this case)

OpenStudy (anonymous):

now all you need to do is solve for x

OpenStudy (anonymous):

let me know if you need to graph these as well

OpenStudy (anonymous):

ok so x it less than or equal to 1

OpenStudy (anonymous):

thank you choprak

OpenStudy (anonymous):

you should get 2 answers: \[x \le -1\] and \[x \ge \frac{ 7 }{ 3 }\]

OpenStudy (anonymous):

anytime you see absolute value (the 2 lines) you should always get 2 answers :P